(b). Use the chain rule to find and so, where w = x2 + y2 + z2, x = st, y = scost, z = s sint when s= 1 and t=0.
Answer #1

2 y=s cost ó zessint (b) given W= x + y² + 2 and u=st ; We have to find ow and do when s=1 and t=0 ds Solution: - W = W(2,4,ot aw ot aw Now calculating aw 2x(3) +246-s.sint) +2z (scost) 2[xs–ys sint + 2560$t] putting x=0; Y=1; 2=0 and 5=1; t=0 2[0 0